The APOcrypha

Volume 81, Issue 8 Saturday, December 10, 1994

Contents

The Editor Takes A Leap Out Of Office 1

Incoming Editorial Comments 1

Outgoing Officer Articles 2 President 2 SVP 3 MVP 5 AVP 7 Treasurer 9 Fellowship Chair 10 APOcEd 11

Pre-Incoming Officer Articles Comment (President) 11

Incoming Officer Articles 12 President 12 SVP 13 Fellowship Rambling #1 14 Treasurer 15

Mold Report IV 17

Fellowship Rambling #2 19

Quotes discuss meeting taken over by APO brothers, news at 11 19

Ye Generic APOc Article 20 Cookie Offer 21 Random Babbling from Gilbert 21 Just a Thought (or a few) 22 pftgoto from bridget 24 Christmas Stories with Kevin 24 Henning’s Random Ramblings “Good Times” 1+1=11 26 Project Report: National Braille Press 28 Project Report: Pledge Project 28 On nice breaks before finals 28 Maria’s Pledge Article 29 Avi’s article 30 Looking Back 31

The Editor Takes A Leap Out Of Office

No, I’m not jumping from the APO office. I’m leaving the office whose duties I’m currently discharging. (I still love that word. It reminds me of nice, romantic things. Like capacitors.) It’s over, I guess. Wrap this up a bit, photocopy, distribute and it’s all Barbara’s. Sniff. I’ll miss it, but you haven’t gotten rid of me yet…

This is, by very very far =), the thickest APOcrypha this term. Maybe Barbara’s better at making you submit then I am. In any case, the fact she’s starting to sign mail with “APOcrypha Dominatrix” isn’t the only reason why I think the APOc is headed for an interesting term...

bert

Incoming Editorial Comments

Watch out! It’s an APOc! =)

I won’t take up pages and pages spewing about how wonderful it to be the editor of the apocrypha. “Why not,” you ask? Because everyone was wonderful and submitted lots and lots of articles, so this edition of the APOcrypha is more than 30 pages long! yay! Thanks to everyone, particularly Gilbert, Henning, Bridget, Karen, Dave, and Chris, who each wrote more than one article, even though Dave gets 200 pushups for turning his in a day late.

That being said, I want to encourage everyone to SUBMIT! to the APOcrypha, as soon as all your finals are over and you have time. Write an article! It doesn’t matter that the next edition of the apocrypha is many weeks away. I am serious about this; since you have time now and during IAP, write a bunch of articles on random things about the chapter, random things about yourself, or just random things in general. Then, when you are hosed with thirty-two problem sets due yesterday and I come begging and pleading for an article, you will have one ready. Better yet, send in the article as soon as you write it, and specify that the next time I even think about trying to bug you for an article, I should just include it and save my bits…

And who knows, if there is enough interest, I may produce a Christmas edition of the apocrypha. Not likely, but possible. =)

Barbara

Outgoing Officer Articles

President

I’m taking the opportunity of our computer system here at ceci failing miserably to write my report. Lukcily my athena connection is still coming through loud and clear.

As president, I can announce quite authoratively that the term happened and the chapter is still alive (and kicking?) We had leadership development (remember barbara’s advice on how to pc a project ?), we had fellowship (remember the candles at chapter weekend II ?), we had service (I still remeber trying to paint up on that roof …), we had wonderful pledges … what more could you ask for ?

Leadership development was helped along by presentations at chapter meetings on ’How to PC a day project’ and ’How to Food PC a project’. More imaginative titles and even different subject matter could easily be thought of for next term. I also tried to instill in the officers a sense of concern for the well being of their PCs - that PCs have to have enough knowledge, experience and training to be able do a good job and need the opportunity to get all of them. Looking to the future, the leadership workshop hosted jointly with BU looks all set for February 11th (mark you calendars NOW) and has generated a lot of enthusiasm already among brothers who want to attend. Possibly, we may then have more ideas about how to design our own leadership development session which could be run more often and at shorter notice.

Inter chapter relations, which while not overtly present, are still being fostered, particularly with the chapter at BU. They will probably take a long time to forget that I gave them 30 service pins, instead of 30 pledge pins, but then it’s taken me all term to live down the embarassment too :) More recently I was at BU’s initiation ceremony (which, funny enough, is almost exactly the same as ours ...) and we went for dinner/ice-cream afterwards. I had a good time, maybe I can convince other brothers to come with me next time. Of course we did schedule joint execcomm to clash with their ceremony, and their semi-formal conflicts with ours; but if we ever get our schedules out of sync ? While I remember, they have invited us to help out at the Cub Scout Decathlon again which they think is on April 1st. (Charley: this is in case I forget to tell you.) Sparky (that EZ alum that’s being showing up a lot recently) has proved a great help in keeping the informal communication between our chapters going. Only 4 brothers went to conclave - but good roll calls come in small packages - and we enjoyed our sojurn in Maine. We have given you all a reputation to keep up at WPI in the Spring (March 10-12, another date for you calendars).

On the national fraternity front, we paid our AAMDS (annual active membership dues) with plenty of time to spare. We found time at chapter meetings and during a dinner trip to discuss a lot of the proposed changes to APO legislation that will be discussed at Nationals. Good luck to dave and ping as they try to figure out what it all means and then map that to the chapter’s opinion polls.

No big strides or achievements were made on two of the chapter’s most consistent, reoccuring problems - communications and membership retention. I have few (no) pearls of wisdom to pass on ... except that more communication is better than less and interdeparmental mail still sucks. Recognising that these areas are in need of constant attention, even though there seems nothing concrete that can be done, is much better than ignoring the issues.

Overall, the term seems to have left everyone with very little time for anything but academics. Hopefully, everyone will have time to recover before the spring term starts and classes can only get easier, right ??

YiLFS Kate

SVP

(courtesy of 16.622, 17.471, 6.003, and 9.00)

In the history of US national security policy... oh wait, wrong article.

I need to be brief, so I encourage everyone and anyone who has questions for me to hunt me down and ask. I plan to stay somewhat actively doing SVP type stuff over IAP and beyond, so I’m very available. Um, I mean... well, you know what I mean.

What went wrong this term:

Not Keeping the PCs Happy In trying to plan ahead for the term, I sometimes lost sight of the problems of the current week. As a result I was slow to respond to (when I even noticed) problems with the project of the week.

Not Achieving all my Lofty Goals I had Big Ideas (tm) of how to make the SVP’s job easier; I also had No Time (tm) to implement them. Well, at least not all of them. But they still exist; I’ll be attempting to implement them over IAP. Chief among them is to revamp the SVP files: check them, update them, and generate an online database of the places we go. I am thinking of something like the general brotherhood listings; something that can be printed to stickers and stuck on cards to provide a physical reference.

Not Finding more People Projects Well, I’m not a huge fan of people projects. Sorry, folks. If you feel really badly about this, please, please, please let the next SVP know.

Of course, some things went well, too:

Projects Happened Yay. =) And in many cases, they happened with a minimum of my effort. Notable among these were Welmet (thanks Gilbert and Andy!), the VA Hospital (thanks Jennifer!), and Chapter Weekend II (thanks Charley!). These projects were found and arranged by the people PC’ing them, which made my job much easier. Future SVPs should note, however, that occasionally checking on things generally makes the PC and the chapter much happier.

Find a Project Study Break Despite some minor glitches (hablas espanol, richard? neither do I…), this was a great way to make the first round of phone calls to sites and enjoy some fellowship as well. Next time, get Bridget to provide cheesecake. =)

New Contacts We worked with some people we have never worked with before, such as the VA hospital. Others have been contacted and are eager to have us come next term. It’s nice to have some new places to go.

There’s much more to say, but if I don’t get some food soon, I may have to eat my words (REALLY BAD PUN ALERT! READ AT OWN RISK! –ed, which is me anyway, but…)[Naah. Seen worse. –ed., which is not me]. And of course I am hosed. But in closing, I would just like to say that being SVP has been one of the more rewarding positions I have held in APO. I really liked the job, and future SVPs looking for help should look me up.

YiLFS, Barbara

MVP

Well. The term is almost over. The pledge term is over, now that voting has started. And here I am, staring at a screen, trying to figure out what to write. Well, here’s my final ExecComm report (Fall 1994 Overview):

What Went Right

• Getting a big brother network up and running was good.

• Trying to get a minimal number of pledge ceremonies was good.

• Rewriting the pledge manual was a big win.

• Giving the manual to the pledges in stages was good.

• Insisting on a pledge service project AND a pledge fellowship event was also good.

• Having a BB meeting at the very beginning of the term was good.

What Went Wrong

• Not enough fellowship events – no one knows anyone else.

• Not finding anyone to help me be in personal contact was bad. Not having time to do it myself was bad.

What Is Ambivalent

• Leaving “the spirit of APO” up to the brothers is chancy. Sometimes it works; sometimes it doesn’t.

• Pledge meetings. I couldn’t think of anything good to have in them that would be commonly appealing. I think not having them in the office might be a good idea.

Suggestions For Next MVP

• Start rush early. Poster like a maniac.

• You can never give prospectives too much information. Trust me.

• Kiss your life goodbye between the start of rush and the pledge ceremony. Read: get help, and lots of it. (Presidents should note that they will want to focus a tremendous amount of their energy on this.)

• Personal (i.e., one-on-one) contact with pledges is vastly preferable to group meetings. However, no one has time for this.

• Being a MVP is not hard. It’s just work. And you have to be a good rolemodel. Everything else can be done with assistance or can be delegated away.

One question I’ve been asking myself at various points during the term is: am I being a good MVP? I don’t know what the answer is. I’ve come to realize that I treat being MVP a lot like being Fellowship Chair – it’s an incredibly nebulous job. Maybe it’s not nebulous; maybe I just don’t grasp what’s necessary. But I find that being MVP is made up of three roles: administrivia, lecturing, and imparting “spirit of APO” to the pledges. The first is routine. The second is difficult to do and not either bore or annoy the pledges. The third is just flat-out impossible to describe How To Do.

It’s also a little scary to realize that the pledge program is Your Department, because as many people feel (myself included), in many ways, the pledge program is the most important facet of the chapter. Reason: no service program for a term results in lack of precedence. However, lack of pledges means a generation gap, which is unfillable. But this is largely irrelevant for me – but highly relevant for Emily. So don’t forget; if the MVP says “jump”, do what she says.

That’s all for now.

dave Outgoing MVP Fall 1994

AVP

I’m off the hook! No more minutes taking! No more ASA meetings! No more room reservations.

Hold on… That sounds too much like Kate’s AVP outgoing report last term. Let me rewind and redo this…

It wasn’t that bad. There are certain things that went wrong though as I pointed them out at the last ExecComm meeting. Let me list them here and discuss them as we go along:

What went wrong:

• I have messed up a couple of the officers’ reports due to my limited minutes-taking ability but, hey, you didn’t send me written reports!

• I had been a little sluggish in printing out up-to-date signup sheets and address lists but we have lots now.

• I skipped the October issue of WUAPO

• I shouldn’t have flamed some officers but officers should correct my mistakes in my “pre-WUAPO” so that we won’t confuse people about the event times and dates in WUAPO.

• I didn’t update the techinfo info about APO yet but I’ll some time after my finals.

What went right:

• Wrote WUAPO generator so that future AVP’s won’t have to keep editing LaTEX and ASCII instances of WUAPO (I still have complete the postscript calendar part.)

• Rooms got reserved.

• Database was updated and modified to indicate visibility on signup sheet so that now the signup sheet is a lot shorter and includes only people who might show up at events. Also featured is new and improved universal perl script.

• Minutes got taken and LaTEX’ed before next meetings. Let me state clearly that this is purely because I’m anally attracted to nice printouts and it is not necessary that minutes be LaTEX’ed.

(Although I do find LaTEX useful as a formatting tool anyway, besides the nice fonts.) If future AVP feels like, they are welcome to use my template and my perl script for converting

.tex file to .txt.

• I made a new on-line minutes browser. The point is not just to impress people with nice formatting, but rather to encourage people to find out about the meetings they missed without roaming through the directory structure of the apo locker. Now you could just do add apo, minutes. This is useful only if people not subscribing to the apo-minutes mailing list know and access the browser. I’ll let the future AVP’s lead this to wherever they want.

• Did some cleaning up for the apo locker, but I’m by no chance done. There’s just too much information in the locker and maintaining it takes a lot of time and also some UNIX proficiency.

Other:

• Henning suggested sending Week at a Glance to apo-news instead of sending mail per event. It is a good idea but a little difficult to implement. From experience, a lot of details of the events don’t get finalized until the last minute and additional mail sent to apo-news may be required to clarify things in addtion to that Week at a Glance. It is, however, possible to “blow up” the current week of the on-line WUAPO to indicate the details of the current events. Of course, you need to let people know about this on-line WUAPO first, and really put an effort into keeping it updated.

• Henning also suggested upgrading the signup sheets to include new columns such as “vege” or “wakeup calls?”. I find this a very good suggestion which Richard has already volunteered to help out. (just a reminder, Richard :).

I think that’s all I can think of for now. (after all, this is not a 40-page thesis :)

Gilbert

Treasurer

It seems somewhat strange to be writing an outgoing officer’s report at this point of the term, but I guess that’s just the feeling the Treasurer has every term. As the responsibility extends until December 31, the end of the fiscal period, and I started with day-to-day stuff end of August, it is hard to give a report that seems somewhat of a final report.

First of all, thank you Kevin for keeping the books up to date while I was gone during the summer, for helping me get acquainted with the job and being available for questions I had.

What has been accomplished (so far)?

• Well, we still have money. That’s always a good thing to hear (or read).

• The Book Exchange account was seperated from the rest of the accounts so that a seperate signature list could be maintained. Furthermore, it is just a good idea to have it seperate from the funds the chapter owns proper.

• As the last of the copied debt list sheets was used up without further copies being made, it became necessary to make up a new form. The latter features an entry for email address (potentially easier to track someone down with in this day and age of “good times”) and is not hand-drawn, but rather was created using a spreadsheet program and looks quite spiffy, if I may say so myself. It resides in the logical space for it in the APO locker:

     /mit/apo/Officers/Treasurer/Forms/debtlist forms.ps

Presented as an example of some of the things that the treasurer sometimes has to deal with is the following ongoing battle, which I mainly included for the amusement value: a billing battle with Carter Rice Dowd, the company that supplies us with drop-poster paper, with the first rounds having been won by their computer, which claims we owe them half a G for an order we never made. The fact that our account name is “MIT, Alpha Phi Omega” makes it the first entry on the list of MIT accounts. It seems it is the default used by a certain salesman. And in that spirit, last week another shipment was delivered to our door. *sigh* Our company rep, Mike (we’re on a first-name basis, we’re old friends, y’know), knows what the problem is and tells me “no problem” and “don’t pay” and “I’ve taken care of it”, yet each month a new statement with more accrued finance charges arrives, giving me another occasion to talk with my pal Mike.

I was really hoping that CPPC (Chapter Program Planning Conference) at Chapter Weekend II was going to be used talking about our financial future, but I was outvoted by the rest of the people at ExecComm who wanted to do a fellowship event and feel good, something admirable as well. Now that we’re feeling good, we can now start thinking about how to guarantee financial survival of the chapter in the future? Therefore, it is planned to get a bunch of interested people together during IAP to mindmeld/brainstorm and make proposals for plans for the future. I see this as being essential to be done soon, i.e. before Registration Day as we know it doesn’t exist anymore.

Pay your debts and I will be grateful!

Henning Colsman-Freyberger

Fellowship Chair

Well, the term is nearly over and we’ve elected new officers to supply fresh energy and ideas to the chapter leadership. It’s time for those of us who were officers this term to step down and let our successors have their shot at changing the way the chapter does things. All the outgoing officers were asked to write outgoing officer reports, so here goes.

I think this term went fairly well from a fellowship point of view. We had all three of the usual major events as well as some study breaks. There weren’t any major disasters and I think everyone had a good time. I didn’t have a master plan when I was elected, I just wanted to make sure there were a reasonable number of study breaks, including past favorites and new ideas. I’m pleased that the bubblewrap study break went over really well. And the people who were at that study break learned that rollerblading over bubblewrap is cool.

I would like to thank all of the people who PC’ed fellowship events this term. The list includes Emily, Jeffrey, Karen, Kate, Ping, and Barbara. If I’ve left anyone’s name out, I apologize. Also, I would like to thank all of the brothers who attended fellowship events and helped make this term’s events a success. Just holding events isn’t all that’s required to have good fellowship in the chapter, though. The brothers have to make an effort to be there for each other, and I believe that’s happened this term. Thanks y’all. Good luck with finals, happy holidays, and have a good IAP.

YiLFS, Rachel

APOcEd

And so my plans to take over the world have been foiled. I’d like to thank all of you that have been writing the articles so I could put the thing together. And to those of you who openly admit to not having ever written an article… we’ll get you yet. And finally, thanks to all of you who are reading this. APOc’s wouldn’t be here without you.

During this term, I’ve had to apologize a lot. My wrists, my lame butt and other factors have been trying to make the APOc late, your articles to get lost, and the locusts… never mind about the locusts. Thank you for your patience.

Editing the APOcrypha is great fun, and I’m sure I’ll miss it. I guess I’ll just have to turn to writing filler or something. =)

I’ll cut this shortnow , lest the APOc overrun its pages… YiLFS, bert

Pre-Incoming Officer Articles Comment (President)

Hi I’m writing an incoming president article as we speak (er, whatever) but first, here’s a word from me as a brother:

theGREisSaturdayAAAIUGH! bridget :)

Incoming Officer Articles

President

Hello all. This will be brief as I am still engaged in reviewing everything I ever learned (or didn’t) in CS.

Although I decided at the start of the term that I should sketch out an updated and more detailed set of plans in case I ran for president again, I didn’t manage to do this, probably because I also tried to front-load the academic year.

So here instead is a rough outline. (You didn’t want to read anything longer, did you?) These are things in which you can help. In fact, without help they will probably sink like rocks.

Communication and information propagation

Some things are getting lost in the current system. As the chapter grows (see membership), loss may increase. I would like people to suggest improvements, or let execcomm know about problems we might not have noticed. Here are a few ideas so far, from various sources, in various stages of being delegated..

1. ‘APO locker tour’ run by an ex-AVP or other clueful person.

2. Signup sheet revision, more general institution of the wakeup call option.

3. A list of the ideal (fastest, most reliable) way to contact each member.

4. Storage of bits of PC clue, somewhere other than in PC’s brain, such as ’the people at XXX take three weeks to return calls’.

Membership: recruiting and retention

The best ’formula’ we have for both of these is Personal Contact. And we know that it’s something that works best when everyone helps. This is difficult for shy people and for busy students, which covers a lot of the chapter. Thanks for your continued efforts :-)

Inactivity

We don’t demand that brothers be active every term; that would be insane considering the sheer effort that academics take. If they’ve been inactive last term, do we tend to assume that they won’t be around this term? Do they still feel welcome around the chapter? In short, is there any factor involved besides academics (over which the chapter has no control). I wonder these things sometimes. (Anyone have some answers, or think the matter is worth talking over?)

Planning

Planning ahead really is useful. To make plans, we have to have goals, ideas, aspirations; otherwise all we do is, at best, maintain the status quo. I’m going to hold a CPPC (chapter program planning conference) in the spring term, at which people in the chapter discuss the direction the chapter is heading and come up with course corrections. It will mostly be useful for next year; if you have any neat ideas for this term, let execcomm or other relevant persons know during IAP if possible. For example, maybe you think it would be keen to invent another big traditional project.. that’s something to talk to the SVP about ideally before the term starts.

Leadership!

It’s nearly midnight and the article will turn into a pumpkin if I don’t mail it soon (hi Barbara) (hi Bridget! -ed). So I’ll just put in a plug for the Leadership Development Workshop, Feb.11. If you don’t consider yourself an old and clueful person, you may well benefit from going (and it’s fun too.) There’ll probably be a size limit so be warned.

One final word. This is your chapter. You, all of us together, make it work. Thanks for trusting me at the wheel, and don’t be afraid to tell us your thoughts and ideas.

YiLFS, Bridget

SVP

Hi, I’m the new SVP elect. My master plan is to unite the world as a single great chapter of APO. My backup plan is to improve communication about projects and to set up better interaction between the SVP and the Fellowship Chair. Wheee!!!!!!

YiLFS and FINALS<sigh>,

Charley

Fellowship Rambling #1

Hi! I’m Karen. Some of you may not know me very well although I’ve been a brother since Spring 1993 (I pledged Fall 1992, my first term at MIT.) I was historian Fall 1993 and have been around in a general sort of way, kinda. Well, SOME people see me. :o)

As fellowship chair next term, I hope WE (as a chapter) can accomplish a number of things as far as fellowship is concerned. But the following list is what I plan to do. I will:

1. Be sure all major Fellowship events like Semi-Formal Formal, Steak Fry, and I-banquet occur as smoothly as possible. (Be kind and volunteer for PCships NOW! Volunteer early, and often.) :oD

2. Make sure study breaks occur throughout the term so brothers can relax a little (punt a little, etc.) and get to know other brothers better. (Sign up to PC these too! Be creative and design your very own.)

3. Encourage brothers who haven’t been as active in the past to come to events once in a while. APO isn’t just a service organization; it’s a service fraterity, and it’s important to let everybody know they are welcome to come to a project even if it is only once in a while. Once you’re a brother, brotherhood is unconditional.

4. I’d like to make the APO office a center for fellowship when there isn’t any organized fellowship stuff. The office is a good place to eat lunch, take a short nap, do a little homework, relax ON-campus, etc. Some brothers already take full advantage of the office, but it would be great to get more brothers using the office. You can even leave things in the office instead of having to go home and pick something up (providing you put your name on it clearly - otherwise you might see it at the next Steak Fry auction.)

5. Work with the SVP and MVP to coordinate events.

6. Do any other warm and fuzzy things that seem appropriate for a Fellowship chair to do. I’m open to ideas. I’m very accessible

     by email. [kgreig@MIT.EDU]

This agenda may be ambitious for ONE person to successfully accomplish, especially since I will be doing other things this term like pitching on the varsity softball team, schoolwork (Course 3), MIT Christian Science Organization and church, Burton deskworking, Concert Band (maybe, hopefully), and other random stuff. But I thought that if I wrote it all down and let brothers think about them, many of these things will naturally occur. Things that YOU as a brother can do to help (besides volunteering to PC stuff): If you have a little (or a big brother or children), be they inactive, active, hyperactive or otherwise, talk to them early next term or over IAP to stay in touch and see what they are doing. Maybe PC and event together. Keeping in touch is a GOOD THING. To many, this is easy - others it is a little more difficult, but it never hurts to care. It is nice to have people check up on you once in a while - good feeling - brother-like. :o)

I hope I can do a good job this coming term. Hope I-banquet goes well, too. (I suppose we all know by now.)

YilFs, Karen Greig :o)

Treasurer

Well, I met with my predecessor, and we came up with a budget for next term. Sometime after classes are over, and we’re both less busy (much, much less busy), we’re going to go over everything else.

My main project this term is going be looking for ways to trim back the budget and otherwise prepare for the ever increasingly likely possibility of Reg Day going away. I’ve already heard a few suggestions: getting brothers to pay for transportation, raising dues, etc., but right now, I’m still very much in the “investigation” phase. I’ll be Proposed Spring 1995 Budget

Prepared by Harry Tsai

Item  Expense Revenue

Service Funds:

CPR  180.00  -
Tools  100.00  -
Transportation  200.00  -
Service Publicity  125.00  -
Service Food (1st Free Meals)  50.00  -
Service Misc.  100.00  -
Executive Discretionary  150.00  -
Membership  180.00  -
Fellowship  200.00  -
Food (Paid)  900.00  900.00
Office (ASC)  100.00  -
Steak Fry Auction  -  100.00
Phone  400.00  20.00
Soda  400.00  400.00
Publicity / WUAPO  20.00  -
Poster (Pubdir)  400.00  400.00
Silkscreen  200.00  150.00
Historian  120.00  -
APOcrypha  150.00  150.00
Registration Day  -  2400.00
National Layaway  300.00  -
Miscellaneous  500.00  400.00
Total Service Funds  4775.00  4920.00

Press Fund:

Press Fund  500.00  500.00
Total Press Fund  500.00  500.00

National Layaway:

National Layaway  -  300.00
Total National Layaway Fund  -  300.00
Grand Total  5275.00  5720.00

going through the books (in minute detail, of course :)) and soliciting advice. We might also form a committee this term if there’s enough interest.

But that’s all a little ways off in the future. Right now, (or a little while ago, depending on when you happen to get to this article,) you all get to discuss and vote on the proposed budget for next term. So… here it is!

Mold Report IV

Greetings dear APOers, most of whom I haven’t seen since the days of UMOC (this being my fault, of course). Hopefully you’re neither too stressed nor too sleepless, but no, I have not yet managed to repress the anguishing memories of end-of-term hecticness (hecticity?) and therefore I do know that non-stress and sleepfulness are not to be found upon/within/among/whatever- other-preposition-you-like you and that thus this is truly a holiday wish very strangely disguised as a finals wish. (And if you understood that sentence, congratulations, because I doubt that I did.)

Anyway, I guess this could be categorized as a “rest-of-you” report that Barbara requested. I predict/promise much randomness, as otherwise it would not be Characteristic Of An Article By Me. Here I go…

Welcome to the world, my new cousin Kevin! (He was born in Houston, over the Thanksgiving holidays.) :)

I saw Star Trek: Generations the night before it premiered (one of those sneak preview thingies). Believe it or not, it was only the second Trek movie I’ve ever seen, and the first (Star Trek IV) I only saw a week before that. A friend of mine rented it for us to watch after a conversation in which she exclaimed (with an appropriate amount of incrdulity, of course), “You graduated from MIT and you’ve NEVER seen a Star Trek movie???????” Heck, I’ve never seen an entire episode of the old Trek. Really! I suppose some/a fanatic few/a frighteningly large number of you think I must be some kind of heathen, but I did enjoy both movies very much, so I can’t be totally hopeless, right? :)

Congratulations to all the new brothers and new officers! Congratulations to all the outgoing officers who made it through this term! Congratulations to anyone who’s becoming mold soon! Congratulations to my sister Jamy who recently made Texas All-State Orchestra! Con-gra-tu-LA-tions, Pro-FES-sor HIG-gins… (And congratulations to anyone who knows where I got that; it’s stuck in my head right now.)

I played three hours of volleyball last Saturday. I actually served six points overhand on a men’s net! Unfortunately, I had game point twice and choked (squawked, too) both times, but I’m getting the hang of the overhand serve. It’s that whole-hand-contact, snap-wrist, elbow-first “throw like a guy” thing. (If you want to know, Just Ask.) It really works!

Okay, last thing. (Whew.) You might see me at I-Banquet. If you do:

Questions NOT to ask me (you’ll risk sounding like my parents):

1. How is the job search going?

2. Are you going to grad school next fall?

3. What have you been doing with all your free time?

4. How’s the money situation?

Questions I will happily answer:

1. Do you have any good books to recommend?

2. How do I buy/join Technique?

3. How do you like your apartment and housemates?

4. Who’s/How’s/Where’s Bob? :)

And that’s all, folks… have a restful break and a very merry IAP!

YiLFS, Amy

Fellowship Rambling #2

Aren’t you all proud of me. This is my second article in ONE APOc and the third one I’ve ever written! (I’m proud! -ed) (The other one was my pledge article - I’ve just been lame about writing articles, I guess.) This one should be titled “Reflections of a Fellowship Chair.” Here goes.

So, what is the relationshop between leadership, fellowship, and service in APO? We separate them into 3 parts, but they are really so intertwined that we must remember that there can’t be one without the other two.

What is service but to be a friend? Fundamentally, they are about the same. We merely differentiate between being of service to others (outside of APO - our service program) and being of service to each other (which is what the fellowship program is all about.)

Where does leadership fit in? Well, someone needs to take the initiative and be of service FIRST - either without waiting to be asked or being ready and willing WHEN asked. Being a leader does not mean taking the initiative and persuading others to friends to you first. Having others of service to you is a consequence, not a starting point. This is the difference between being a good leader and a bad politician.

Always keep in mind that APO comprises all three of these things equally. It’s like the U.S. government. The checks and balances are in place, inherent in the system, because it was designed that way.

Yilfs, Karen :o)

Quotes discuss meeting taken over by APO brothers, news at 11

[4962] sorokin@ATHENA.MIT.EDU quotes 12/01/94 18:01
Subject: autumn
Matt and I have decided that we are ready to make the
committment of sharing a SIPB drawer. We understand the
responsibilites this entails, and if you would like to

donate my drawer to someone else, you may. =)
[4963] marc@ATHENA.MIT.EDU quotes 12/01/94 22:14
Subject: dcctdw
I have style, and I have a big mouth, and that's it.
[4964] rjbarbal@ATHENA.MIT.EDU quotes 12/02/94 21:03
Subject: ckclark
Sunsoft's biggest problem is that they are not the Borg,
but they fantasize that they are.  For example, they are
under the delusion that people will actually run Solaris
on Intel processors, and that they will take over this
way.  In reality, there are about 4 people in the world
using 2.4 on Intel machines, they all work at Sun, and
they boot Windows when the boss isn't looking.

(courtesy of Ping)

Ye Generic APOc Article

Well, ok. It’s not terribly generic.

It’s a lichen report. Next term, i’ll be a lichen, in all probability. What that means is that i’ll be a part-time student, finishing my thesis, but hopefully working part-time. So, I’ll probably go inactive next term, because I still feel it’s The Right Thing To Do – all that stuff about getting out of the way of the actives.

So maybe I should go to AXAA meetings. Oooh, what a scary concept.

It’s kinda interesting – I don’t know what I want to do with APO now that I’ve basically graduated/escaped. (It’s unclear if i’ll graduate. It is clear that I’ve almost escaped.) I’ve given some thought to infiltrating AXAA and bending its considerable power to my will. (Author’s note: HA!) (Oh, what the heck, i’ll throw this in for free – Editor’s note: HA! (Hey! Who’s the editor here, anyway? -ed)) I’ve thought about serving on staff. But I think there’s something to be said for just holding a job, paying off debts, living life, and mellowing out from the rush of undergraduate life. Hmm. I guess what I’m saying is that I don’t feel mature enough to be on staff.

But for next term, I’ll be around MIT, being involved in Ballroom Dance Team and (cough) hopefully other things as well. :) I have no clue what i’ll be doing in 2 years, but working in Boston for the next year seems a reasonable guess.

dave

Cookie Offer

Some of you may have noticed that the only part of the office which really rivals the messiness of Dave’s desk is the one covered with ancient photographs of various molds and fungi. As historian, I’ve determined that these no longer recognizable specimens are still of historical interest and need to be better preserved. And I really want to leave a clean desk.

So, I am hereby announcing the 1995 First and Hopefully Last Annual IAP Cleanup of Ancient Photographs. This will take place the weekend before IAP starts (January 6-9). Since I have no clue who in the world most of the people in the photographs are, I need the help of other ancient molds and… er, I mean, the help of the Alumni and other more senior members of the Chapter. So, if you will be in the area just before IAP starts, watch your email (I’ll send to apo-news@mit.edu) and the office bulletin board for times I’ll be going through cool pictures of the Chapter’s past.

This should be a really cool fellowship thing, and since I will not be hosed, I will very likely make some yummy chocolate chip cookies.

+ Richard Historian and Cookie-Baker

Random Babbling from Gilbert

I really doubt if many people know why I am hosed this term. Let me use the opportunity to explain: I am going out with too many girls: Dina (Dynamic Systems, 6.241), Ana (Analysis, Complex, 18.04), Stocky (Stochastic Processes Detection and Estimation, 6.342), Maggie (Macroeconomics, 14.02), Melody (Music Composition, 21.M351), Toby (Tau Beta Pi), April (APO), and Klavia (Piano). Some of these ladies can only be pleased by Matlab scripts rather than flowers or love letters. Some likes sensitive touches and agile fingers… In any case, I am dumping Ana and Stocky because the former is too boring and the latter thinks that I’m too stupid (I’ll chase after her again next term because I like challenging things. Playing hard-to-get, huh? :)) I would suggest to everyone not to get too ambitious and go out with lots of guys and girls. Stick to 5 or 6 and that’s should be enough to get you by the term without feeling lonely.

Of course, being a random babbling, it wouldn’t just end it this soon. I have announcements to make. I’m earnestly arranging the weddings (concerts) with both Melody and Klavia (boy, how stressful!). (multiple weddings are ok. Just make sure that they don’t conflict with each other.) You are all invited:

Time  Place  Concert  Comment

Thu 12/15, 2pm Killian Hall Compositions from 21.M351 (solos and duets)

Our new brother Grant is one of the performers!

Sat 12/17, 3pm Killian Hall Piano recital by students of Marek Zebrowski I’ll be playing a Debussy prelude and first movement of a Mozart sonata

While I would love to have your presence, I do want to warn you that the former has mostly atonal music. Take this as you will…

Gilbert

Just a Thought (or a few)

I plan to become inactive next term for various reasons (One is that I would like to concentrate my energy on other, maybe new, things as my MIT life might be coming to an end.) I just have some comments/suggestions (that I remember. I think I might have more that I forgot…) to make as this might be my last issue of Apoc while I’m still active.

Open and Close I always thought that the “open” in open meetings and open projects sound pretty fascist. It implies that the other meetings and projects we have are “closed” to outsiders while we do welcome prospectives who don’t even plan to have any affiliation with apo whatsoever to go to these events and explore APO. Couldn’t we just change them to “Info Meetings” and “Service Projects” and make it clear that people plan to pledge are required to go to 1 of the first X Info meetings and 1 of the first Y service projects?

Membership Recruitment I might be too sensitive but it seems that the percentage of pledges getting initiated seems to have dropped. And more are on the borderline case in terms of making the requirements. Are we pushing people too hard into pledging? Are we not letting prospectives know what the pledgeship entails? I think it helps the chapter’s morale to concentrate on bringing the pledges up to brotherhood. Dave’s big brother meeting was a good one but it should probably be done a little earlier in the term and emphasize on the solutions to problems.

Membership Retention/Fellowship Dave sharply pointed out that there is a difference between “fellowship” and “social”. We are in fact losing a number of brothers. While it is true that they could be hosed, I’m not convinced that they don’t spend any time with their friends. if they would feel comfortable enough about the chapter and are friends with the brothers, they would hang out around APO anyway. The problem is that we are probably not very close to achieving this. We could take a low-key approach and let whoever gets along with the chapter be around. But if we have a goal of expanding the chapter, we should put an effort into becoming friends with the people different from us, people who might be harder to know because they are quieter, etc. I heard that the fellowship event at Chapter Weekend II went really well: people sat in a circle and talked about their experience (mostly positive) with APO. This is a great start as it reinforces people’s stay at APO. Except that it didn’t include the brothers who are already inactive. Membership retention is a hard problem about which I have little ideas. But, for each inactive brother/pledge, which I’ll call X for the sake of clarity, there’s a brother Y whom X feels the most comfortable with. This might be X’s big brother or someone from X’s pledge class, etc. If we find out who this is, it makes a big difference for X if Y keeps in touch with X and encourages X to attend APO events. I personally had positive experience with this and think that this is something worth pursuing or at least thinking about.

Social event idea As an individual who always strives for something new, I think we could run something like a “Movie Marathon” during IAP. Movie rentals are only $1.xx at Towers from Tue-Thu. The chapter has a VCR already and I’m happy to lend the chapter my ancient yet fully functional 19” TV for the event. All the PC needs to do is to take a survey of what people like to see, announce the event, and rent the 5-6 movies from Towers, set up the TV/VCR in the office (maybe move the furniture a little bit). W20 is such a strategic location that if someone doesn’t like a certain movie, he/she could always have other places (not limited to Athena) to hang out and come back for another one. Food as usual is a necessity :)

Gilbert

pftgoto from bridget

I have a new fish tank.. you can drop by Senior House and say hi to the fish after I-banquet if you want :-) -bridget

Christmas Stories with Kevin

Hi everyone!

Well, its almost Christmas! Hey, I have a question for all my fellow Christians. Do you have Jewish friends? (I do.) Have you always felt like a butthead because you never get around to wishing your Jewish friends “Happy Haunakah” until Christmas, and even then you can’t spell it? (I have.) Did you swear to yourself that this would be the year you finally do that on the actual first day of Haunakah? (I did.) And did you blow it anyway? (I did. Am I weird or does everyone do this?)

Anyway, since its Christmas I’m sending in my favorite Christmas story. Its all about me and my friend Ed, who was APO president and my apartment mate back in our WPI days. Since we had 5 APO brothers in the apartment, we had agreed to host the annual APO Holiday party.

Ed and I were sitting around our apartment late one night noticing that we hadn’t decorated our apartment for Christmas. We also noticed that we had no money for Christmas decorations, and even if we had money we didn’t have much time to shop that week (most of our free time ocurred between the hours of midnight and 8 am, when few malls are open) and even if we had time we both hated shopping.

So this is what we did.

We got a stack of index cards and a box of magic markers. Ed wrote “Wreath” on one of the index cards and hung it on the outside door of our apartment. I wrote “Candle” on an index card and stuck it in the window. This happy exchange continued and the rest of our apartment mates got in on the act until our apartment was decorated with “Mistletoe,” “Bowl of candy canes,” “Cardboard cut out of Santa Claus,” “Mistletoe,” “Christmas tree,” “Mistletoe,” “String of popcorn” hanging from the tree, “More mistletoe,” several “Ornaments” also hanging from the tree, “Mistletoe Mistletoe Mistletoe”, “Fireplace,” “Good grief, more mistletoe!” and of course, “Ed’s stocking” and “Kevin’s stocking” which were hanging on the wall above the fireplace. (Ed at that time was dating a charming young lady named Linda, who is now his wife, but the rest of the residents of that apartment were pretty desperate dudes, as the above list implies.) I was also able to locate a brick, so I wrote “Fruitcake” on it and stuck it on the kitchen table. We also wrote “Broken ornament” on an index card, ripped it in half and placed it on the floor under the tree.

The party was very successful. A good time was had by all. Everyone got to make a stocking to put over the fireplace, and it you wanted to give someone a present, you just wrote it into their stocking. Someone gave me a queen of spades. If you play hearts you understand how nasty this was. I also received “A lump of coal,” “A boot to the head” and “A pound of salt.” (Bear in mind, my nickname in the chapter was Slug.)

Ah yes, life was always fun at WPI. I really enjoyed being the among the chapter’s favorite whipping posts for verbal abuse. Those were the good old days.

Hrm. Come to think of it, these new days really aren’t all that different…

Henning’s Random Ramblings “Good Times” 1+1=11

Last column I asked the following:

Who is the highest, who the lowest-paid governor? (Which states?)

Two people sent in a guess to trivia@mit.edu. Gilbert Leung (gratuitous mention #1) was first to send something and suggested California and Nebraska. Pete Wilson made $120,000 this year, but he’s still “only” the fourth-highest net gubernatiorial grosser. The Lincoln office is not quite the lowest-paid gubernatiorial office with $65 K. Surpassing it in lower pay are only Bill Clinton’s former office ($60,000) and Marc Racicot’s Montanan office, the lowest-yielding one, which pays $55,502 yearly. The latter was identified by Jeff Bigler (gratutious mention #2), who also correctly determined the highest-paid governor to come from my home state New York, Mario Cuomo and his $130,000 salary.

Kevin Dahm (gm #3) answered with a simple, to-the-point “Who cares? They all get paid too damn much. Or am I getting them mixed up with professional athletes?” Well, trust Kevin to bring baseball into a discussion. :) I guess every one should decide for themselves whom they derive more utility from, Roger Clemens or Governor Weld…

For the record, the seven remaining seven gratuitous mentions: Ping, Kate, Rachel, Gilbert (he sent two things, thus mentioned again), Alexis, Alexis and Alexis (that’s how I interpret “me, me, mention me” :) And also how I get to ten mentions.).

During IAP, I am running a few activities of which two might be of interest to some of you: Eighties Movie Nights (Activity #142) and NY Movie Nights (#154), running Tuesday and Thursday nights respectively. Such great movies such as Ferris Bueller’s Day Off, Weird Science, Cadillac Man, The Freshman, The Taking of Pelham One Two Three, Say Anything, Ghostbusters will be shown. If you are interested in being on the mailing lists announcing the finalized schedules, drop a line to 80s-films-request and/or ny-films-request.

The following appeared in the International Herald Tribune in 1919, 75 years ago on December 8:

Fancy schools

LONDON - Great controversy is raging over the anti-snobbery campaign of the great public schools of England [Henning’s note: UK public school = US private school]. Eton and Harrow seem the chief victims, because smartness in dress, including the proverbial top-hat, signify snobbery to the �plebs� who are incapable of judging the difference between manners and re�nement on the one hand and conceipt and pose on the other.

Quite. :) They don’t write ’em as they used to.

Two questions to keep you thinking (or not) until the next publication:

1. When did the NYSE Dow Jones index first close above 1000? (FYI, it’s around 3750 now.)

2. What year did the MIT Lightweight Crew team win the Thames Challenge Cup at Henley, England? (Hint: The answer can be found on the inbound platform of the Kendall Square T stop.)

If you know one or both answers, email it to trivia@mit.edu!

And yet another term’s end quickly approaches. Good luck on finals and enjoy your winter break.

(There was a slight error in the URL mentioned last week. Here is the new, shorter URL:

http://www.mit.edu:8001/people/henning/trivia.html)

Project Report: National Braille Press

The project was ran by Barbara and myself. Overall I thought that the project went very well. However, one item that I was disappointed with was brothers who said they would show and did not come or call us or let us know. This project was interesting because it was the same day as Project America, the nation-wide community service day. This meant that we got free breakfast as we were waiting to listen to Dan Quayle speak. He was late so we decided to go to project. Ping is reported as saying, “Why don’t we go up to the microphone and announce that we wish we could stay, but we are going to go do project now.” After working to move piles of zinc, stacking music and painting a bathroom not big enough for two cats—much less Bridget and a couple of other people. We were then treated to Uno’s and continued working until about 2:00pm.

Christopher

Project Report: Pledge Project

This project went well. However, more organizanization was needed… I believe that greater cooperation is needed between the MVP and SVP and pledge PC or PCs. Also, on a side note, the pledge PC should have a key to the office for a Saturday project, in particular if a borother requested a 6am wake-up call.

Christopher

On nice breaks before finals

Well, after my announcement of being done until finals at the end of Joint ExecComm I was instantly the recipient of flames from a hostile brotherhood. However, the threat over e-mail was most effective (thank you Big Bro). Well, here it is.

Question: How do I earn a nice break before finals?

Answer: don’t take labs, Unifried, three 00*’s, 1.00 etc. Those late lab reports and problem sets can be a bummer. Finish work ahead of the time, at the end there is nothing to do. Just punt the problem sets, not a problem if you are a frosh— however, upperclasmen should consider the effects on their GPA. Take 8.022, 7.012, 18.023, 24.00. 18.023 has no problem sets. The last 7.012 pset was due on Dec 1. The last 8.022 set was due Dec 7. The last paper for 24.00 was due Dec 4. This is the secret for having nothing due in the week before finals…

Well, I am finished spewing, and I felt that Neophyting went well and thank you for not barging into The Tech.

Christopher

P.S. Pledge article note: I understand there was a discussion of the word “icshe”. It is Cherokee for “crap”.

Maria’s Pledge Article

[This got eaten by my mailbox, but I bravely rescued it… –bert]

My name is Maria Pritykin. I got it a long time ago, I’d suppose some time on February 22nd, 1976.

I’ve decided to be a part of APO on the reg day when Jeffrey was asking people to come help out with the Book Exchange. I really wanted to.

Well, now a little bit about myself and my hobbies. I’ve liked to draw and paint since I remember myself. With some help from different people, I started getting involved in mathematics. Well, I declared my major: 18 (applied), but I’m still interested in art-related courses. My plans for the future: become a better person. No, this is not meant as a cliché, but as a goal that is hard to reach. Some things that worry me are what I can do to reduce stress in my life because too much of it resulted in just too many problems. Well, back to my plans. I’d like to do what I plan to and plan reasonably. I’ll start with “sleep.”

Avi’s article

In response to bert’s requests for brothers to write articles about themselves, perhaps it is alos appropriate for alumni to do the same— even if I might seem to even predate mold…

I pledged APO twenty-seven years ago. I was first introduced to AX at a summer rush held at Mike Bromberg’s for incoming freshmen. The chapter did a good job of entertaining us and answering our questions, doing a very soft-sell on APO. I picked up some literature before getting a ride home with Al Stiehl.

At MIT, Bob Boyd stopped by my room at Baker to see if I was going to the open rush, as I had expressed interest that summer. I responed that I was ineligible to join APO, having never been in Scouting. Bob revealed to me that the chapter had a liberal interpretation of that factor. The president, Al Brenstein, had never been a Scout. AX felt that joining APO, which was nationally affiliated to BSA, made you able to belong to APO. A circuitous logic! Anyway, during my term as a pledge, MVP Ron Rosen attended a constitutional convention in Norman, OK, and led a successful effort to change the pledging requirement to following Scout ideals instead of being a member. When I became a Brother, I met the letter of the by-laws, not just the intent.

The members of the James A. Lovell Pledge Class didn’t write APOcrypha articles, but we did list a paragraph of interests. Bern (Thor) Krafsig included Peanuts in his list. The capital P caught my attention. Somehow, the friendship that developed still runs strong, as it has with other members of the pledge class and the chapter. I also remember the pledge quiz. The first question was to name the founder of APO. I knew his first name was Frank and the last started with an H. Going to Inquisition (the formal interview held back then), I was sure that I would be a continuing pledge. For some reason, the chapter took me in on my weak pledge performance.

Over my four years as an undergraduate, APO was the most important factor in my life. I was very active in the service program, held a few offices, developed strong, lasting friendships, and perhaps learned more from and about APO than in my classes. Some said I may have earned my degree in VII, but I really majored in APO.

The summer after graduating, I even met my wife on an APO outing. APO unquestionably did something to me, as I still have a strong interest in it.

This only scratches the surface, but it can be a form of introduction and can open topics for discussion when I’m up at MIT. I look forward to getting to know more of you in the future. (I’m sorry that I was unable to make it to last month’s Steak Fry & Auction.)

The Blue Dragon (aka Avi Ornstein)

Looking Back

Having just read some of the pledge articles, and having reflected over the years in writing an article about myself, I felt that pledging might be an appropriate topic for this (infrequent) column.

The Blue Dragon

… there were 17 pledges at the first smoker, 9 at the second. Eleven have signed up. For the past two years, the fall pledge class has numbered between 35 and 40 at its inception.

Carl Mampaey, MVP (F’62) [XXII:4]

[Note: we now use “prospectives” where Carl used “pledges”. –Avi]

Today’s membership report has been cancelled due to the impending holiday [Thanksgiving]. Suffoce it to say that neither the last nor the next pledge meetings have been cancelled.

Abba Weinstein, MVP (F’63) [XXIII:4]

… Uncle Dave will shortly be sending you your private list of freshmen to contact, help and cherish. Answer any questions they ahve and leave your name, address, and phone number for their further use. Be friendly, helpful, don’t push the chapter too much. Our best advertising is guys who are interested in serving others.

You might also try to hold a rush party with the other brothers in the area… Make it a ‘come and talk about MIT and APO’ session combined with a lot of fun and fellowship, where you fill them up with food, joy, and info about the ’tute and AX.

Len Tower, SVP (F’69) [XXX:1]

We have a myriad of pledges (20 < myriad < 30)

Chapter meeting minutes (10/14/70) [XXXII:5]

An accepted tradition (for many, many years now) is that each pledge class gets to throw their MVP into some body of water after Spring Afterfinals.

Avi Ornstein (F’71) [XXXIV:3]

Among so many people, one feels a lesser part of the whole. Pledges are being raised not knowing their fellow peons… Members lose the feeling of togetherness as the Fellowship Circle grows.

Sheva Stern, letter (S’72) [XXXV:3]

Membership VP (Richards): … I’d rather have fewer pledges with more attention per pledge, for a lower inactive rate.

ExecComm minutes (9/6/72) [XXXVI:3]

As in the past terms, pledges will be encouraged to submit a short introductory write-up of themselves.

Chapter Meeting minutes (9/27/72) [XXXVI:4]

The APOcrypha

Volume 81, Issue 8

Fall 1994 ExecComm

President  Kate Curtis  628-5384 kcurtis
SVP  Barbara Manganis x5-7408  autumn
MVP  Dave Cho  497-2189 dcctdw
AVP  Gilbert Leung  x5-7241  gleung

Treasurer Henning Colsman-

Freyberger  x5-7659  henning
Fellowship  Rachel Caileff  x5-7414  rcaileff
APOcEd  bert Dvornik  576-2773 bert
ASC  Cathy Coury  734-9211 cmcoury
PubDir  Charley Hamilton x5-8941  chashmlt

Historian Richard Barbalace 497-2189 rjbarbal Chair, Advisory Comm. Joe Kuchta

Brotherhood Expectations

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 5 Chapter Meetings and 1 ExecComm Learn, teach, or practice a chapter skill Hold a leadership position within the chapter (e.g. PC, sub-PC, hold office) Attend the Pledge Term Review Meeting Attend ceremonies

Pledge Requirements

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 2 Chapter Meetings and 1 ExecComm Attend at least two-thirds of scheduled Pledge Meetings Learn or teach a chapter skill Get a Big Brother within two weeks of pledging PC or sub-PC a service project or fellowship event Attend the Pledge Feedback Meeting